Data for Hermann Corporation are shown below:

Per Unit Percent of Sales
Selling price $ 70 100 %
Variable expenses 49 70
Contribution margin $ 21 30 %

Fixed expenses are $74,000 per month and the company is selling 4,400 units per month.

2-a. Refer to the original data. How much will net operating income increase (decrease) per month if the company uses higher-quality components that increase the variable expense by $4 per unit and increase unit sales by 25%.

2-b. Should the higher-quality components be used?
Answer #1
Current Proposed
Unit Sales 4400 (4400*1.25)=5500
Sales (4400*70)=308,000 (5500*70)=385,000
Variable expenses (4400*49)=215600 5500*(49+4)=291500
Contribution margin 92400 93500
Fixed expenses 74,000 74,000
Net operating income 18400 19500

Increase in Net operating income=(19500-18400)=$1100

Hence higher quality component should be used.